a plane flying horizontally at 350 ms^-1 releases a package at an altitude of 1.5 * 10^3 m calculate:
1. how long does it take the package to reach the ground
2. how far will it move horizontally while falling
1
Expert's answer
2021-10-08T14:59:22-0400

1)

"h=\dfrac{1}{2}gt^2,""t=\sqrt{\dfrac{2h}{g}}=\sqrt{\dfrac{2\cdot1.5\cdot10^3\ m}{9.8\ \dfrac{m}{s^2}}}=17.5\ s."

2)

"x=v_{0x}t=350\ \dfrac{m}{s}\cdot17.5\ s=6125\ m."